### v4

# Sine laws on semigroups with an involutive anti-automorphism: A Levi–Civita approach via left translations

Abstract.
Stetkær’s matrix (Levi–Civita) method is a powerful tool for functional equations on semigroups involving a homomorphism $\sigma$, as it yields a finite-dimensional invariant space under right translations and a corresponding matrix formalism. However, this framework collapses when $\sigma$ is an involutive anti-automorphism due to the order reversal in the right-regular action. In this paper, we overcome this obstruction at the operator level by establishing the conjugation identity: letting $J$ denote composition with $\sigma$, we prove
$J\,R(\sigma(y))\,J=L(y)\qquad(\forall\,y\in S),$
which converts the problematic right translates into left translations. Using this left-translation approach, we obtain an anti-automorphic Levi–Civita closure principle and apply it to the generalized sine law, recovering the classical dichotomy $\beta\in\{\pm 1\}$ and the associated $xy$-addition law. Moreover, under a mild centrality assumption (e.g. for commutative semigroups), we also recover the standard $\sigma$-transformation rules (such as $f\circ\sigma=\beta f$ and the corresponding formula for $g\circ\sigma$).
Key words and phrases:
Functional equation, Sine addition law, Semigroup, Anti-automorphism, Levi-Civita equation, Central function.
2020 Mathematics Subject Classification:
39B52; 20M15.
1. Introduction
The classical trigonometric addition laws (most notably d’Alembert’s equation and the sine addition law) form a cornerstone of the theory of functional equations. Since the pioneering monographs (e.g. [1]), a major theme has been to understand how the algebraic structure of the underlying domain controls the analytic structure of solutions. Over the past decades, this program has been pushed far beyond abelian groups to non-abelian groups [5] and, more recently, to semigroups, where the absence of inverses and commutativity forces genuinely new techniques.
A modern semigroup-oriented line of research, developed in particular by Ebanks [3, 4] and Stetkær [6, 7], treats generalized trigonometric-type equations on a semigroup $S$ involving a map $\sigma:S\to S$ (typically a homomorphism, an involutive automorphism, or closely related variants). A central technical tool in this literature is the Levi–Civita method: one shows that a finite-dimensional space generated by solutions is invariant under a suitable translation action, and then exploits that invariance to derive strong structural constraints.
To explain the obstruction addressed in this paper, consider the two-parameter family of generalized sine addition/subtraction laws
$f(x\sigma(y))=f(x)g(y)+\beta\,g(x)f(y)+\gamma\,f(x)f(y),\qquad\beta\in F^{\ast},\ \gamma\in F,$ (1.1)
studied systematically in [6]. When $\sigma$ is a homomorphism, Stetkær’s approach (implemented via right translations) yields a sharp dichotomy and shows that linearly independent solutions satisfy a simpler $xy$-addition law together with precise transformation rules such as $f\circ\sigma=\beta f$.
However, the situation changes when $\sigma$ is an anti-homomorphism, i.e. $\sigma(xy)=\sigma(y)\sigma(x)$. Then the natural right action no longer interacts with the product in the same way:
$R(\sigma(y_{1}))\,R(\sigma(y_{2}))=R(\sigma(y_{1})\sigma(y_{2}))=R(\sigma(y_{2}y_{1})).$ (1.2)
This order reversal prevents a direct transfer of the usual Levi–Civita arguments based on right translates. Such anti-automorphisms arise naturally in many settings, for instance matrix transposition, inversion in groups, and involutions on symmetry structures. A principled semigroup theory in the anti-automorphic setting therefore calls for an adapted translation-operator framework.
Operator-level viewpoint. When $\sigma$ is an involutive anti-automorphism, the standard Levi–Civita method based on right translations breaks down because the family $\{R(\sigma(y))\}_{y\in S}$ satisfies the order-reversing composition law (1.2). We therefore work at the operator level: letting $J$ denote composition with $\sigma$, $(Jh)(x)=h(\sigma(x))$, we prove the conjugation identity
$J\,R(\sigma(y))\,J=L(y)\qquad(y\in S),$ (1.3)
which converts the problematic right translates into a left-translation action on the conjugated space $J(V)$.
Main idea: conjugation to left translations. Throughout, $\sigma:S\to S$ is assumed to be a surjective involutive anti-automorphism. The conjugation identity (1.3) (Lemma 3.1) allows us to transfer the right-translate action $R(\sigma(\cdot))$ to left translations, which restores the translation-invariance mechanism needed for a Levi–Civita analysis. In the application to the generalized sine law, a mild additional centrality hypothesis is only used to obtain the explicit $\sigma$-transformation rule for $g$.
Main results. Our first main result is an anti-automorphic Levi–Civita closure theorem (Theorem 3.2) for equations of the form
$f(x\sigma(y))=f(x)h_{1}(y)+g(x)h_{2}(y),$
with $\{f,g\}$ and $\{h_{1},h_{2}\}$ linearly independent. The theorem shows that $J(V)$ is invariant under left translations (equivalently, $V$ is invariant under $R(\sigma(\cdot))$), yielding a closure principle tailored to involutive anti-automorphisms.
Our second main result applies this closure to the generalized sine law (1.1). In Theorem 4.4 we obtain an analogue of Stetkær’s structural theorem: for linearly independent solutions $(f,g)$, one necessarily has $\beta\in\{\pm 1\}$ and the equation forces a simpler $xy$-addition law. Furthermore, assuming the existence of a central element $c\in Z(S)$ with $f(c)\neq 0$ (for instance, when $S$ is commutative and $f\not\equiv 0$), the solutions satisfy the expected $\sigma$-transformation rules:
$f\circ\sigma=\beta f,\qquad g\circ\sigma=\begin{cases}g+af,&\beta=-1,\\ g,&\beta=1.\end{cases}$
In particular, the anti-automorphic setting does not introduce new exceptional parameter regimes; instead, the conjugation viewpoint recovers the classical dichotomy.
Organization of the paper. Section 2 collects the needed definitions and notation (central/abelian functions, parity, and the regular actions $R$ and $L$). In Section 3 we develop the conjugation-based framework: we prove (1.3) and establish the anti-automorphic Levi–Civita closure theorem (Theorem 3.2). Section 4 applies this closure to the generalized sine law and proves our main structural classification (Theorem 4.4), where an additional mild centrality hypothesis is assumed only to obtain the explicit $\sigma$-transformation rule for $g$. Section 5 provides concrete illustrations from matrix groups and the symmetric group, where the conjugation identity can be seen explicitly. Finally, Section 6 gives concluding remarks and connections to recent centrality/abelianity results in the literature.
2. Preliminaries
Throughout, $S$ is a semigroup with multiplicative notation; $F$ is a field with $\mathrm{char}(F)\neq 2$; $F^{\ast}=F\setminus\{0\}$. We write $F(S,F)$ for the space of $F$-valued functions on $S$.
Definition 2.1 (Central, abelian [7]).
A function $F:S\to A$ is central if $F(xy)=F(yx)$ for all $x,y\in S$. A function $F:S\to F$ is abelian if
$F(x_{1}x_{2}\cdots x_{n})=F(x_{\pi(1)}x_{\pi(2)}\cdots x_{\pi(n)})$
for all $n\geq 2$ and all permutations $\pi$. A central function is abelian if and only if it satisfies Kannappan’s condition $F(xyz)=F(xzy)$ for all $x,y,z\in S$.
Definition 2.2 (Parity [6, 7]).
Given a map $\varphi:S\to S$, a function $F:S\to F$ is called even with respect to $\varphi$ if $F\circ\varphi=F$, and odd with respect to $\varphi$ if $F\circ\varphi=-F$.
Definition 2.3 (Homomorphisms, anti-homomorphisms).
A map $\varphi:S\to S$ is a homomorphism if $\varphi(xy)=\varphi(x)\varphi(y)$, and an anti-homomorphism if $\varphi(xy)=\varphi(y)\varphi(x)$. An involutive anti-automorphism is a bijective anti-homomorphism $\sigma$ with $\sigma^{2}=\mathrm{id}$.
Definition 2.4 (Right/left regular representations).
The operators $R,L:S\to\mathrm{End}\big(F(S,F)\big)$ are defined by
$(R(y)h)(x):=h(xy),\qquad(L(y)h)(x):=h(yx).$
Then $R$ is a (semi)group homomorphism and $L$ is an anti-homomorphism:
$R(y_{1})R(y_{2})=R(y_{1}y_{2}),\qquad L(y_{2})L(y_{1})=L(y_{1}y_{2}).$
3. A Left-Translation Method via Operator Conjugation
Throughout this section, $\sigma:S\to S$ is a surjective anti-automorphism with $\sigma^{2}=\mathrm{id}$.
3.1. Conjugating Right Translates into Left Translates
Define an involutive linear operator $J:F(S,F)\to F(S,F)$ by
$(Jh)(x):=h(\sigma(x)).$
Since $\sigma$ is involutive, $J^{2}=\mathrm{id}$ and $J^{-1}=J$.
Lemma 3.1 (Conjugation Identity).
For all $y\in S$, the operator identity holds:
$J\,R(\sigma(y))\,J=L(y).$
Proof.
Fix $h\in F(S,F)$ and $x\in S$. Using the definitions of $J$ and $R$, and the fact that $\sigma$ is an anti-homomorphism and involutive, we compute
$(JR(\sigma(y))Jh)(x)$ $=(R(\sigma(y))Jh)(\sigma(x))$
$=(Jh)(\sigma(x)\sigma(y))$
$=h\!\left(\sigma(\sigma(x)\sigma(y))\right)$
$=h\!\left(\sigma(\sigma(yx))\right)\qquad(\text{since }\sigma(x)\sigma(y)=\sigma(yx))$
$=h(yx)\qquad(\text{since }\sigma^{2}=\mathrm{id})$
$=(L(y)h)(x).$
3.2. Anti-automorphism Levi–Civita closure
Theorem 3.2.
Let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Suppose $f,g,h_{1},h_{2}:S\to F$ satisfy
$f(x\sigma(y))=f(x)h_{1}(y)+g(x)h_{2}(y),\quad x,y\in S,$ (3.1)
where $\{f,g\}$ and $\{h_{1},h_{2}\}$ are linearly independent. Set $V:=\mathrm{span}\{f,g\}$ and $V^{\sigma}:=J(V)=\mathrm{span}\{Jf,Jg\}$. Then, $V^{\sigma}$ is invariant under $L(y)$ for every $y\in S$. Equivalently, $V$ is invariant under $R(\sigma(y))$ for every $y\in S$.
Proof.
Write (3.1) in operator form:
$R(\sigma(y))f=h_{1}(y)f+h_{2}(y)g.$
Conjugate by $J$ and use Lemma 3.1:
$L(y)(Jf)=JR(\sigma(y))J(Jf)=J(R(\sigma(y))f)=h_{1}(y)Jf+h_{2}(y)Jg.$
Thus $L(y)(Jf)\in V^{\sigma}$.
We now show that $L(y)(Jg)\in V^{\sigma}$ for all $y\in S$. Indeed, choose $y_{0}\in S$ such that $h_{2}(y_{0})\neq 0$ (exists since $h_{1},h_{2}$ are linearly independent). From
$L(y_{0})(Jf)=h_{1}(y_{0})Jf+h_{2}(y_{0})Jg$
we obtain
$Jg=\frac{1}{h_{2}(y_{0})}\,L(y_{0})(Jf)-\frac{h_{1}(y_{0})}{h_{2}(y_{0})}\,Jf.$
Now fix $y\in S$ and apply $L(y)$:
$L(y)(Jg)=\frac{1}{h_{2}(y_{0})}\,L(y)L(y_{0})(Jf)-\frac{h_{1}(y_{0})}{h_{2}(y_{0})}\,L(y)(Jf).$
Since $L$ is an anti-homomorphism (Definition 2.4), $L(y)L(y_{0})=L(y_{0}y)$. Hence
$L(y)(Jg)=\frac{1}{h_{2}(y_{0})}\,L(y_{0}y)(Jf)-\frac{h_{1}(y_{0})}{h_{2}(y_{0})}\,L(y)(Jf).$
We already know $L(z)(Jf)\in V^{\sigma}$ for all $z\in S$, so the right-hand side lies in $V^{\sigma}$. This proves the claim, hence $V^{\sigma}$ is $L$-invariant. ∎
4. The Generalized Sine Law with an Anti-Automorphism
In this section we study the generalized sine law
$f(x\sigma(y))=f(x)g(y)+\beta\,g(x)f(y)+\gamma\,f(x)f(y),\qquad x,y\in S,$ (4.1)
where $\beta\in F^{\ast}$, $\gamma\in F$, and $f,g$ are linearly independent.
Lemma 4.1 (A coefficient dichotomy).
Let $S$ be a semigroup and $F$ a field with $\mathrm{char}(F)\neq 2$. Assume that $f,g:S\to F$ are linearly independent and that
$f(xy)=f(x)g(y)+b\,g(x)f(y)+c\,f(x)f(y)\qquad(\forall x,y\in S),$
where $b,c\in F$ are constants. Then $b\in\{0,1\}$.
Proof.
See [7, Lemma 5.1]; cf. also [6, Lemma 3.3]. ∎
Lemma 4.2.
Assume (3.1). Define
$\widehat{h}_{1}:=h_{1}\circ\sigma,\qquad\widehat{h}_{2}:=h_{2}\circ\sigma.$
Then
$f(xy)=f(x)\widehat{h}_{1}(y)+g(x)\widehat{h}_{2}(y)\qquad(\forall\,x,y\in S).$
Proof.
Replace $y$ by $\sigma(y)$ in (3.1) and use $\sigma^{2}=\mathrm{id}$:
$f(xy)=f\bigl(x\sigma(\sigma(y))\bigr)=f(x)h_{1}(\sigma(y))+g(x)h_{2}(\sigma(y))=f(x)\widehat{h}_{1}(y)+g(x)\widehat{h}_{2}(y).$
Lemma 4.3 (Central-element criterion for $\sigma$-conjugate span).
Let $S$ be a semigroup and let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy
$f(x\sigma(y))=f(x)\,h(y)+u(x)\,f(y)\qquad(\forall x,y\in S),$ (4.2)
where $f$ and $u$ are linearly independent. Assume moreover that
$f\circ\sigma=\varepsilon f\qquad\text{for some }\varepsilon\in\{\pm 1\}.$ (4.3)
If there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$, then
$u\circ\sigma\in\mathrm{span}\{f,u\}.$
In particular, the conclusion holds whenever $S$ is commutative and $f\not\equiv 0$.
Proof.
Fix $y\in S$ and $x\in S$. Since $\sigma$ is an involutive anti-automorphism,
$\sigma(yx)=\sigma(x)\sigma(y)\quad\Rightarrow\quad yx=\sigma(\sigma(x)\sigma(y)).$
Using (4.3), we have
$(L(y)f)(x)=f(yx)=f\!\left(\sigma(\sigma(x)\sigma(y))\right)=\varepsilon\,f(\sigma(x)\sigma(y)).$
Now apply (4.2) to $(x,y)=(\sigma(x),y)$:
$f(\sigma(x)\sigma(y))=f(\sigma(x))h(y)+u(\sigma(x))f(y).$
Using $f(\sigma(x))=\varepsilon f(x)$ and $u(\sigma(x))=(u\circ\sigma)(x)$, we obtain
$f(\sigma(x)\sigma(y))=\varepsilon f(x)h(y)+(u\circ\sigma)(x)\,f(y).$
Multiplying by $\varepsilon$ and using $\varepsilon^{2}=1$, we get the identity of functions
$L(y)f=h(y)\,f+\varepsilon f(y)\,(u\circ\sigma)\qquad(\forall y\in S).$ (4.4)
Now let $c\in Z(S)$ with $f(c)\neq 0$. Since $\sigma$ is surjective, pick $y_{0}\in S$ such that $\sigma(y_{0})=c$. Then
$(R(\sigma(y_{0}))f)(x)=f(x\sigma(y_{0}))=f(xc)=f(cx)=(L(c)f)(x),$
where we used $c\in Z(S)$. Applying Theorem 3.2 to (4.2) shows that $V:=\mathrm{span}\{f,u\}$ is invariant under $R(\sigma(\cdot))$, hence $R(\sigma(y_{0}))f\in V$, so $L(c)f\in V$.
Finally, take $y=c$ in (4.4):
$L(c)f=h(c)\,f+\varepsilon f(c)\,(u\circ\sigma).$
Since $L(c)f\in V$ and $f(c)\neq 0$, we solve for $u\circ\sigma$:
$u\circ\sigma=\frac{1}{\varepsilon f(c)}\,L(c)f-\frac{h(c)}{\varepsilon f(c)}\,f\in V.$
Thus $u\circ\sigma\in\mathrm{span}\{f,u\}$. ∎
Theorem 4.4.
Let $\sigma:S\to S$ be a surjective involutive anti-automorphism, and assume that there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$. Suppose $f,g:S\to F$ are linearly independent and satisfy (4.1).
  1. (i)
    If $\beta=-1$, then $\gamma=0$. Furthermore, there exists $a\in F$ such that
    f(xy)=f(x)g(y)+g(x)f(y)+af(x)f(y),f(xy)=f(x)g(y)+g(x)f(y)+a\,f(x)f(y),
    and
    $f\circ\sigma=-f,\qquad g\circ\sigma=g+af.$
  2. (ii)
    If $\beta\neq-1,$ then necessarily $\beta=1$, and
    $f(xy)=f(x)g(y)+g(x)f(y)+\gamma\,f(x)f(y),$
    and
    $f\circ\sigma=f,\qquad g\circ\sigma=g.$
Proof.
We distinguish the two cases $\beta=-1$ and $\beta\neq-1$.
(i) The case $\beta=-1$. Then (4.1) becomes
$f(x\sigma(y))=f(x)\bigl(g(y)+\gamma f(y)\bigr)-g(x)f(y)\qquad(x,y\in S).$ (4.5)
Set
$h:=g+\gamma f,\qquad u:=-g.$
Then (4.5) can be written in the Levi–Civita form (3.1) as
$f(x\sigma(y))=f(x)\,h(y)+u(x)\,f(y)\qquad(x,y\in S),$ (4.6)
i.e. (3.1) holds with $(f,g,h_{1},h_{2})=(f,u,h,f)$. By Lemma 4.2 we therefore have, for all $x,y\in S$,
$f(xy)=f(x)h(\sigma(y))+u(x)f(\sigma(y))=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y)).$ (4.7)
Step 1: a two–variable law and parity of $f$. Apply Theorem 3.2 to (4.6). Then
$V:={\rm span}\{f,u\}={\rm span}\{f,g\}$
is invariant under $R(\sigma(z))$ for every $z\in S$. Hence, for each $z\in S$, there exist functions $P,Q:S\to F$ such that
$g(x\sigma(z))=P(z)\,f(x)+Q(z)\,g(x)\qquad(\forall x\in S).$ (4.8)
Fix $x,y,z\in S$. Since $\sigma$ is an anti-automorphism, $\sigma(yz)=\sigma(z)\sigma(y)$, so $f(x\sigma(yz))=f(x\sigma(z)\sigma(y))$. Compute both sides using (4.5).
(1) With $yz$:
$f(x\sigma(yz))=f(x)\bigl(g(yz)+\gamma f(yz)\bigr)-g(x)f(yz).$
(2) With $x$ replaced by $x\sigma(z)$:
$f(x\sigma(z)\sigma(y))=f(x\sigma(z))\bigl(g(y)+\gamma f(y)\bigr)-g(x\sigma(z))f(y).$
Insert (4.5) for $f(x\sigma(z))$ and (4.8) for $g(x\sigma(z))$, then compare coefficients of the linearly independent functions $f(x)$ and $g(x)$. This yields, for all $y,z\in S$,
$f(yz)=f(z)g(y)+Q(z)f(y)+\gamma f(z)f(y).$ (4.9)
Putting $y=x$ gives
$f(xz)=f(z)g(x)+Q(z)f(x)+\gamma f(z)f(x)\qquad(\forall x,z\in S).$ (4.10)
On the other hand, taking $y=\sigma(z)$ in (4.5) gives
$f(xz)=f(x)g(\sigma(z))-g(x)f(\sigma(z))+\gamma f(x)f(\sigma(z))\qquad(\forall x,z\in S).$ (4.11)
Comparing (4.10) and (4.11) and using linear independence of $f$ and $g$ yields, for all $z\in S$,
$f(\sigma(z))=-f(z),\qquad g(\sigma(z))=Q(z)+2\gamma f(z).$ (4.12)
In particular, $f\circ\sigma=-f$.
Step 2: obtain $g\circ\sigma\in\mathrm{span}\{f,g\}$. Rewrite (4.6) as
$f(x\sigma(y))=f(x)\,h(y)+u(x)\,f(y),\qquad h=g+\gamma f,\ \ u=-g.$
From Step 1 we already have $f\circ\sigma=-f$, i.e. (4.3) holds with $\varepsilon=-1$. Applying Lemma 4.3 to (4.2), we conclude that
$u\circ\sigma\in\mathrm{span}\{f,u\}.$
Equivalently,
$g\circ\sigma\in\mathrm{span}\{f,g\}.$ (4.13)
Step 3: exclude $g\circ\sigma=-g+\mu f$. By (4.13) there exist $\lambda,\mu\in F$ such that
$g\circ\sigma=\lambda g+\mu f.$
Apply $\sigma$ again and use $f\circ\sigma=-f$:
$g=\lambda(g\circ\sigma)+\mu(f\circ\sigma)=\lambda(\lambda g+\mu f)-\mu f=\lambda^{2}g+(\lambda\mu-\mu)f.$
By linear independence of $f$ and $g$, we obtain
$\lambda^{2}=1,\qquad(\lambda-1)\mu=0.$
Assume $\lambda=-1$. Then $g\circ\sigma=-g+\mu f$. Replacing $y$ by $\sigma(y)$ in (4.5) and using $f\circ\sigma=-f$ gives
$f(xy)=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))=f(x)\bigl(-g(y)+\mu f(y)-\gamma f(y)\bigr)+g(x)f(y).$
Let $\widehat{g}:=-g$. Then
$f(xy)=f(x)\widehat{g}(y)-\widehat{g}(x)f(y)+(\mu-\gamma)\,f(x)f(y),$
which contradicts Lemma 4.1. Hence $\lambda\neq-1$, so $\lambda=1$ and
$g\circ\sigma=g+af\quad\text{for some }a\in F.$ (4.14)
Step 4: derive $\gamma=0$ and finish. From (4.7), using $f\circ\sigma=-f$ and (4.14), we obtain for all $x,y\in S$,
$f(xy)$ $=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))$
$=f(x)\bigl((g(y)+af(y))+\gamma(-f(y))\bigr)-g(x)(-f(y))$
$=f(x)\bigl(g(y)+(a-\gamma)f(y)\bigr)+g(x)f(y).$
Thus
$f(xy)=f(x)g(y)+g(x)f(y)+(a-\gamma)f(x)f(y)\qquad(\forall x,y\in S).$ (4.15)
Now compute f(xy)-f(xy) in two ways. From (4.15),
$f(xy)=-f(x)g(y)-g(x)f(y)-(a-\gamma)f(x)f(y).$
On the other hand, since $f\circ\sigma=-f$ and $\sigma$ is an anti-automorphism,
$f(xy)=(f\circ\sigma)(xy)=f(\sigma(xy))=f(\sigma(y)\sigma(x)).$
Apply (4.5) with $(x,y)=(\sigma(y),x)$ and use $f\circ\sigma=-f$ and (4.14):
$-f(xy)$ $=f(\sigma(y)\sigma(x))=f(\sigma(y))\bigl(g(x)+\gamma f(x)\bigr)-g(\sigma(y))f(x)$
$=(-f(y))g(x)+(-f(y))\gamma f(x)-\bigl(g(y)+af(y)\bigr)f(x)$
$=-f(y)g(x)-g(y)f(x)-(a+\gamma)f(x)f(y).$
Comparing the two expressions for f(xy)-f(xy) yields $\gamma=0$. With $\gamma=0$, (4.15) becomes
f(xy)=f(x)g(y)+g(x)f(y)+af(x)f(y),f(xy)=f(x)g(y)+g(x)f(y)+a\,f(x)f(y),
and we already have $f\circ\sigma=-f$ and $g\circ\sigma=g+af$. This proves (i).
(ii) The case $\beta\neq-1$. Define
$g_{1}:=g+\frac{\gamma}{1+\beta}f.$
Then $f$ and $g_{1}$ are linearly independent and (4.1) is equivalent to
$f(x\sigma(y))=f(x)g_{1}(y)+\beta\,g_{1}(x)f(y)\qquad(x,y\in S).$ (4.16)
Apply Theorem 3.2 to (4.16). Then $V:={\rm span}\{f,g_{1}\}$ is invariant under $R(\sigma(z))$ for every $z\in S$. Hence, for each $z\in S$, there exist functions $P,Q:S\to F$ such that
$g_{1}(x\sigma(z))=P(z)\,f(x)+Q(z)\,g_{1}(x)\qquad(\forall x\in S).$ (4.17)
Using associativity as before, $f(x\sigma(yz))=f(x\sigma(z)\sigma(y))$. Computing both sides via (4.16) and using (4.17), and then comparing coefficients of the linearly independent functions $f(x)$ and $g_{1}(x)$, we obtain for all $y,z\in S$,
$g_{1}(yz)$ $=g_{1}(z)g_{1}(y)+\beta\,P(z)f(y),$ (4.18)
$f(yz)$ $=f(z)g_{1}(y)+Q(z)f(y).$ (4.19)
Putting $y=x$ in (4.19) gives
$f(xz)=f(z)\,g_{1}(x)+Q(z)\,f(x)\qquad(\forall x,z\in S).$ (4.20)
Taking $y=\sigma(z)$ in (4.16) gives
$f(xz)=f(x)\,g_{1}(\sigma(z))+\beta\,g_{1}(x)f(\sigma(z))\qquad(\forall x,z\in S).$ (4.21)
Comparing (4.20) and (4.21) yields, for all $z\in S$,
$Q(z)=g_{1}(\sigma(z)),\qquad f(z)=\beta\,f(\sigma(z)).$ (4.22)
Applying $\sigma$ to $f(z)=\beta f(\sigma(z))$ and using $\sigma^{2}={\rm id}$ gives $\beta^{2}=1$. Since $\beta\neq-1$, we conclude
$\beta=1\quad\text{and hence}\quad f\circ\sigma=f.$ (4.23)
Step 2: show $g_{1}\circ\sigma=g_{1}$. From Step 1, we already have $\beta=1$ and $f\circ\sigma=f$. With $\beta=1$, equation (4.16) reads
$f(x\sigma(y))=f(x)g_{1}(y)+g_{1}(x)f(y)\qquad(\forall x,y\in S).$
This is exactly (4.2) with $u=h=g_{1}$ and $\varepsilon=1$ in (4.3). Applying Lemma 4.3, we get
$g_{1}\circ\sigma\in\mathrm{span}\{f,g_{1}\}.$
Thus $g_{1}\circ\sigma=\lambda g_{1}+\mu f$ for some $\lambda,\mu\in F$. Applying $\sigma$ again and using $f\circ\sigma=f$ yields
$g_{1}=\lambda(g_{1}\circ\sigma)+\mu(f\circ\sigma)=\lambda(\lambda g_{1}+\mu f)+\mu f=\lambda^{2}g_{1}+(\lambda\mu+\mu)f.$
Hence $\lambda^{2}=1$ and $(\lambda+1)\mu=0$. If $\lambda=-1$, then $g_{1}\circ\sigma=-g_{1}+\mu f$, and replacing $y$ by $\sigma(y)$ in (4.16) (with $\beta=1$) gives
$f(xy)$ $=f(x)g_{1}(\sigma(y))+g_{1}(x)f(\sigma(y))$
$=f(x)\bigl(-g_{1}(y)+\mu f(y)\bigr)+g_{1}(x)f(y)$
$=-f(x)g_{1}(y)+g_{1}(x)f(y)+\mu f(x)f(y).$
Let $\widehat{g}_{1}:=-g_{1}$. Then
$f(xy)=f(x)\widehat{g}_{1}(y)-\widehat{g}_{1}(x)f(y)+\mu f(x)f(y),$
contradicting Lemma 4.1. Therefore $\lambda\neq-1$, so $\lambda=1$ and (since $\mathrm{char}(F)\neq 2$) we must have $\mu=0$. Consequently,
$g_{1}\circ\sigma=g_{1}.$ (4.24)
Finally, apply Lemma 4.2 to (4.16) with $\beta=1$, and use (4.23) and (4.24). We obtain for all $x,y\in S$,
$f(xy)=f(x)g_{1}(\sigma(y))+g_{1}(x)f(\sigma(y))=f(x)g_{1}(y)+g_{1}(x)f(y).$
Substituting $g_{1}=g+\frac{\gamma}{2}f$ gives
$f(xy)=f(x)g(y)+g(x)f(y)+\gamma f(x)f(y)\qquad(\forall x,y\in S).$
Moreover, from $g_{1}\circ\sigma=g_{1}$ and $f\circ\sigma=f$ we get
$g\circ\sigma+\frac{\gamma}{2}f=g_{1}\circ\sigma=g_{1}=g+\frac{\gamma}{2}f,$
hence $g\circ\sigma=g$. This proves (ii). ∎
5. Examples: Concrete Illustrations of the Conjugation Identity
We illustrate the conjugation identity on familiar groups (hence semigroups). In each example, $\sigma$ is an involutive anti-automorphism and $J$ denotes composition with $\sigma$,
$(Jh)(x)=h(\sigma(x)).$
We verify the operator identity of Lemma 3.1:
$J\,R(\sigma(y))\,J=L(y)\qquad(\forall\,y\in S).$
For the chosen finite-dimensional space $V$ of coordinate functions, we also describe explicitly how the left-translation operators $L(y)$ act on $V$.
We emphasize that these examples are included solely to illustrate the conjugation identity $J\,R(\sigma(y))\,J=L(y)$ and the resulting left-translation action; they are independent of the additional centrality assumption used later in Theorem 4.4.
5.1. Example 1: $GL_{n}(K)$ with $\sigma(A)=A^{\top}$ and column-coordinate functions
Let $K$ be a field and let $S=GL_{n}(K)$ with $\sigma(A)=A^{\top}$. Then $\sigma$ is an involutive anti-automorphism since $(AB)^{\top}=B^{\top}A^{\top}$ and $(A^{\top})^{\top}=A$. Fix $v\in K^{n}$ with $v\neq 0$ and define, for $i=1,\dots,n$,
$f_{i}(A):=e_{i}^{\top}A\,v=(Av)_{i},\qquad A\in GL_{n}(K),$
where $(e_{1},\dots,e_{n})$ is the standard basis of $K^{n}$. Put $V:=\mathrm{span}\{f_{1},\dots,f_{n}\}$.
Verification of $JR(\sigma(y))J=L(y)$. Here $(Jf_{i})(A)=f_{i}(A^{\top})=e_{i}^{\top}A^{\top}v$. For $y\in GL_{n}(K)$ and $X\in GL_{n}(K)$,
$\big(JR(\sigma(y))Jf_{i}\big)(X)$ $=\big(R(\sigma(y))Jf_{i}\big)(\sigma(X))=\big(R(y^{\top})Jf_{i}\big)(X^{\top})$
$=(Jf_{i})(X^{\top}y^{\top})=e_{i}^{\top}(X^{\top}y^{\top})^{\top}v=e_{i}^{\top}(yX)v$
$=f_{i}(yX)=(L(y)f_{i})(X).$
Hence $JR(\sigma(y))J=L(y)$.
Left action on $V$. For $y\in GL_{n}(K)$ and $X\in GL_{n}(K)$,
$(L(y)f_{i})(X)=f_{i}(yX)=e_{i}^{\top}yXv=\sum_{j=1}^{n}y_{ij}\,e_{j}^{\top}Xv=\sum_{j=1}^{n}y_{ij}\,f_{j}(X).$
In particular, $L(y)V\subseteq V$ for every $y\in GL_{n}(K)$.
5.2. Example 2: The symmetric group $S_{n}$ with $\sigma(\pi)=\pi^{-1}$
Let $S=S_{n}$ and $\sigma(\pi)=\pi^{-1}$. Fix $j_{0}\in\{1,\dots,n\}$ and define, for $i=1,\dots,n$,
$f_{i}(\pi):=\mathbf{1}_{\{\pi(j_{0})=i\}},\qquad\pi\in S_{n},$
and set $V:=\mathrm{span}\{f_{1},\dots,f_{n}\}$.
Verification of $JR(\sigma(y))J=L(y)$. Here $(Jf_{i})(\pi)=f_{i}(\pi^{-1})=\mathbf{1}_{\{\pi(i)=j_{0}\}}$. For $y\in S_{n}$ and $x\in S_{n}$,
$\big(JR(\sigma(y))Jf_{i}\big)(x)$ $=\big(R(\sigma(y))Jf_{i}\big)(\sigma(x))=\big(R(y^{-1})Jf_{i}\big)(x^{-1})$
$=(Jf_{i})(x^{-1}y^{-1})=(Jf_{i})\big((yx)^{-1}\big)=f_{i}(yx)=(L(y)f_{i})(x).$
Hence $JR(\sigma(y))J=L(y)$.
Left action on $V$. For $y\in S_{n}$ and $x\in S_{n}$,
$(L(y)f_{i})(x)=\mathbf{1}_{\{(yx)(j_{0})=i\}}=\mathbf{1}_{\{x(j_{0})=y^{-1}(i)\}}=f_{y^{-1}(i)}(x).$
Thus $L(y)$ permutes the spanning set $\{f_{1},\dots,f_{n}\}$, and in particular $L(y)V\subseteq V$.
5.3. Example 3: $SO(3,\mathbb{R})$ with $\sigma(g)=g^{-1}=g^{\top}$ and column functions
Let $S=SO(3,\mathbb{R})$ and $\sigma(g)=g^{-1}=g^{\top}$. Write $g=(g_{ij})_{1\leq i,j\leq 3}$ and define
$f_{1}(g):=g_{11},\qquad f_{2}(g):=g_{21},\qquad f_{3}(g):=g_{31},\qquad V:=\mathrm{span}\{f_{1},f_{2},f_{3}\}.$
Verification of $JR(\sigma(y))J=L(y)$. We have $(Jf_{i})(X)=f_{i}(X^{\top})=X_{1i}$. For $y\in SO(3,\mathbb{R})$ and $x\in SO(3,\mathbb{R})$,
$\big(JR(\sigma(y))Jf_{i}\big)(x)$ $=\big(R(\sigma(y))Jf_{i}\big)(\sigma(x))=\big(R(y^{\top})Jf_{i}\big)(x^{\top})$
$=(Jf_{i})(x^{\top}y^{\top})=f_{i}\big((x^{\top}y^{\top})^{\top}\big)=f_{i}(yx)=(L(y)f_{i})(x).$
Hence $JR(\sigma(y))J=L(y)$.
Left action on $V$. Since $(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,x_{j1}$, we get for $y,x\in SO(3,\mathbb{R})$,
$(L(y)f_{i})(x)=f_{i}(yx)=(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,f_{j}(x).$
In particular, $L(y)V\subseteq V$.
Remark 5.1 (Linear independence in Examples 1–3).
The spanning families used above are linearly independent.
Example 1. Assume $v\neq 0$. If $\sum_{i=1}^{n}c_{i}f_{i}\equiv 0$, then for all $A\in GL_{n}(K)$,
$0=\sum_{i=1}^{n}c_{i}f_{i}(A)=\Big(\sum_{i=1}^{n}c_{i}e_{i}^{\top}\Big)Av=c^{\top}(Av).$
Since $GL_{n}(K)$ acts transitively on $K^{n}\setminus\{0\}$, we have $\{Av:A\in GL_{n}(K)\}=K^{n}\setminus\{0\}$, hence $c=0$.
Example 2. If $\sum_{i=1}^{n}c_{i}f_{i}\equiv 0$, choose $\pi$ with $\pi(j_{0})=k$ to get $c_{k}=0$.
Example 3. If $c_{1}f_{1}+c_{2}f_{2}+c_{3}f_{3}\equiv 0$, evaluate at $I$ to get $c_{1}=0$. Choose rotations sending $e_{1}$ to $e_{2}$ and $e_{1}$ to $e_{3}$ to get $c_{2}=c_{3}=0$.
The examples above show concretely how Lemma 3.1 converts the family $R(\sigma(y))$ into left translations $L(y)$ via conjugation by $J$, and how this mechanism can be checked explicitly on natural finite-dimensional spaces of coordinate functions.
6. Concluding Remarks
The conjugation identity $J\,R(\sigma(y))\,J=L(y)$ provides the key link needed to transfer the anti-automorphic right-translate action into left translations, thereby avoiding the order-reversal obstruction in the family $R(\sigma(\cdot))$. This left-translation viewpoint yields a clean Levi–Civita closure principle in the anti-automorphic setting and leads to analogues of Stetkær’s structural results for the generalized sine law, including the dichotomy $\beta\in\{\pm 1\}$ and the induced $\sigma$-parity rule for $f$. Moreover, under a mild additional centrality hypothesis (e.g. for commutative semigroups), one also recovers the full $\sigma$-transformation rule for $g$. The examples illustrate the conjugation mechanism concretely on standard semigroups and show how the relevant finite-dimensional spaces of coordinate functions behave under left translations.
Funding.
None.
Availability of data and materials
Not applicable.
Declarations
Conflict of interest
None.

### v5

@@
 which converts the problematic right translates into left translations.
-Using this left-translation approach, we obtain an anti-automorphic Levi–Civita closure principle and apply it to the generalized sine law, recovering the classical dichotomy $\beta\in\{\pm 1\}$ and the associated $xy$-addition law.
-Moreover, under a mild centrality assumption (e.g.
-for commutative semigroups), we also recover the standard $\sigma$-transformation rules (such as $f\circ\sigma=\beta f$ and the corresponding formula for $g\circ\sigma$).
+Using this left-translation approach, we obtain an anti-automorphic Levi–Civita closure principle and apply it to the generalized sine law.
+Remarkably, the classical dichotomy $\beta\in\{\pm 1\}$ and the parity relation $f\circ\sigma=\beta f$ are recovered unconditionally.
+Furthermore, under a natural bridge hypothesis, which is automatically satisfied when there exists a central element $c$ with $f(c)\neq 0$, we obtain the corresponding standard $xy$-addition law and the exact $\sigma$-transformation rule for $g$.
 Key words and phrases:
-Functional equation, Sine addition law, Semigroup, Anti-automorphism, Levi-Civita equation, Central function.
+Functional equation, Sine addition law, Semigroup, Anti-automorphism, Levi–Civita equation, Translation invariance.
 2020 Mathematics Subject Classification:
@@
 A central technical tool in this literature is the Levi–Civita method: one shows that a finite-dimensional space generated by solutions is invariant under a suitable translation action, and then exploits that invariance to derive strong structural constraints.
-To explain the obstruction addressed in this paper, consider the two-parameter family of generalized sine addition/subtraction laws
+To motivate the operator viewpoint adopted here, consider the two-parameter family of generalized sine addition/subtraction laws
 $f(x\sigma(y))=f(x)g(y)+\beta\,g(x)f(y)+\gamma\,f(x)f(y),\qquad\beta\in F^{\ast},\ \gamma\in F,$ (1.1)
@@
 When $\sigma$ is a homomorphism, Stetkær’s approach (implemented via right translations) yields a sharp dichotomy and shows that linearly independent solutions satisfy a simpler $xy$-addition law together with precise transformation rules such as $f\circ\sigma=\beta f$.
-However, the situation changes when $\sigma$ is an anti-homomorphism, i.e. $\sigma(xy)=\sigma(y)\sigma(x)$. Then the natural right action no longer interacts with the product in the same way:
+When $\sigma$ is an anti-homomorphism, i.e. $\sigma(xy)=\sigma(y)\sigma(x)$, the parametrized family $y\mapsto R(\sigma(y))$ satisfies the order-reversing composition law
 $R(\sigma(y_{1}))\,R(\sigma(y_{2}))=R(\sigma(y_{1})\sigma(y_{2}))=R(\sigma(y_{2}y_{1})).$ (1.2)
-This order reversal prevents a direct transfer of the usual Levi–Civita arguments based on right translates.
-Such anti-automorphisms arise naturally in many settings, for instance matrix transposition, inversion in groups, and involutions on symmetry structures.
-A principled semigroup theory in the anti-automorphic setting therefore calls for an adapted translation-operator framework.
+Thus the family $R(\sigma(\cdot))$ is no longer a right-regular representation in the parameter $y$.
+Of course, because $\sigma^{2}=\mathrm{id}$, one may rewrite (1.1) by replacing $y$ with $\sigma(y)$ and thereby recover a standard $xy$-equation.
+Our aim here is complementary: we isolate an intrinsic operator identity explaining how the anti-automorphic family $R(\sigma(\cdot))$ is converted into a genuine translation action.
 Operator-level viewpoint.
-When $\sigma$ is an involutive anti-automorphism, the standard Levi–Civita method based on right translations breaks down because the family $\{R(\sigma(y))\}_{y\in S}$ satisfies the order-reversing composition law (1.2).
-We therefore work at the operator level: letting $J$ denote composition with $\sigma$, $(Jh)(x)=h(\sigma(x))$, we prove the conjugation identity
+When $\sigma$ is an involutive anti-automorphism, the family $\{R(\sigma(y))\}_{y\in S}$ is order-reversing in the parameter $y$ by (1.2).
+Letting $J$ denote composition with $\sigma$, $(Jh)(x)=h(\sigma(x))$, we prove the conjugation identity
 $J\,R(\sigma(y))\,J=L(y)\qquad(y\in S),$ (1.3)
-which converts the problematic right translates into a left-translation action on the conjugated space $J(V)$.
+which converts the family $R(\sigma(\cdot))$ into left translations. This gives a natural operator-level reformulation of the anti-automorphic situation and is the viewpoint used throughout the paper.
 Main idea: conjugation to left translations.
-Throughout, $\sigma:S\to S$ is assumed to be a surjective involutive anti-automorphism.
+Throughout, $\sigma:S\to S$ is assumed to be an involutive anti-automorphism.
 The conjugation identity (1.3) (Lemma 3.1) allows us to transfer the right-translate action $R(\sigma(\cdot))$ to left translations, which restores the translation-invariance mechanism needed for a Levi–Civita analysis.
-In the application to the generalized sine law, a mild additional centrality hypothesis is only used to obtain the explicit $\sigma$-transformation rule for $g$.
-Main results. Our first main result is an anti-automorphic Levi–Civita closure theorem (Theorem 3.2) for equations of the form
+In the application to the generalized sine law, the decisive additional input is a bridge condition of the form
+$\exists\,d\in S:\qquad f(d)\neq 0\quad\text{and}\quad L(d)f\in\mathrm{span}\{f,g\}.$
+A central element $c\in Z(S)$ with $f(c)\neq 0$ is a convenient sufficient condition for this bridge assumption, but the bridge formulation is the intrinsic one used in the proof.
+Main results. Our first main result, Theorem 3.2, is an anti-automorphic Levi–Civita closure theorem for equations of the form
 $f(x\sigma(y))=f(x)h_{1}(y)+g(x)h_{2}(y),$
-with $\{f,g\}$ and $\{h_{1},h_{2}\}$ linearly independent.
+where $\{f,g\}$ is linearly independent and $h_{2}\not\equiv 0$.
 The theorem shows that $J(V)$ is invariant under left translations (equivalently, $V$ is invariant under $R(\sigma(\cdot))$), yielding a closure principle tailored to involutive anti-automorphisms.
 Our second main result applies this closure to the generalized sine law (1.1).
-In Theorem 4.4 we obtain an analogue of Stetkær’s structural theorem: for linearly independent solutions $(f,g)$, one necessarily has $\beta\in\{\pm 1\}$ and the equation forces a simpler $xy$-addition law.
-Furthermore, assuming the existence of a central element $c\in Z(S)$ with $f(c)\neq 0$ (for instance, when $S$ is commutative and $f\not\equiv 0$), the solutions satisfy the expected $\sigma$-transformation rules:
+In Theorem 4.6, under the bridge assumption above, we obtain an analogue of Stetkær’s structural theorem: for linearly independent solutions $(f,g)$, one necessarily has $\beta\in\{\pm 1\}$, the equation yields the standard $xy$-addition law, and the expected $\sigma$-transformation rules
 $f\circ\sigma=\beta f,\qquad g\circ\sigma=\begin{cases}g+af,&\beta=-1,\\ g,&\beta=1.\end{cases}$
-In particular, the anti-automorphic setting does not introduce new exceptional parameter regimes; instead, the conjugation viewpoint recovers the classical dichotomy.
+Moreover, the proof shows that the parity relation $f\circ\sigma=\beta f$ is obtained before the bridge assumption is invoked, whereas the bridge assumption is used only to obtain control of $g\circ\sigma$ and thereby to reach the standard $xy$-addition law.
+Corollary 4.7 records the frequently useful central-element criterion, and Example 5.4 shows that the bridge assumption is strictly weaker.
 Organization of the paper.
@@
 In Section 3 we develop the conjugation-based framework: we prove (1.3) and establish the anti-automorphic Levi–Civita closure theorem (Theorem 3.2).
-Section 4 applies this closure to the generalized sine law and proves our main structural classification (Theorem 4.4), where an additional mild centrality hypothesis is assumed only to obtain the explicit $\sigma$-transformation rule for $g$.
-Section 5 provides concrete illustrations from matrix groups and the symmetric group, where the conjugation identity can be seen explicitly.
+Section 4 applies this closure to the generalized sine law and proves our main structural classification under the bridge assumption; Corollary 4.7 gives the corresponding result for central elements.
+Section 5 provides concrete illustrations from matrix groups and the symmetric group, including a non-central example showing that the bridge assumption can hold even when no central element $c$ satisfies $f(c)\neq 0$.
 Finally, Section 6 gives concluding remarks and connections to recent centrality/abelianity results in the literature.
@@
 3. A Left-Translation Method via Operator Conjugation
-Throughout this section, $\sigma:S\to S$ is a surjective anti-automorphism with $\sigma^{2}=\mathrm{id}$.
+Throughout this section, $\sigma:S\to S$ is an involutive anti-automorphism.
 3.1. Conjugating Right Translates into Left Translates
@@
 Theorem 3.2.
-Let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Suppose $f,g,h_{1},h_{2}:S\to F$ satisfy
+Let $\sigma:S\to S$ be an involutive anti-automorphism. Suppose $f,g,h_{1},h_{2}:S\to F$ satisfy
 $f(x\sigma(y))=f(x)h_{1}(y)+g(x)h_{2}(y),\quad x,y\in S,$ (3.1)
-where $\{f,g\}$ and $\{h_{1},h_{2}\}$ are linearly independent.
+where $\{f,g\}$ is linearly independent and $h_{2}\not\equiv 0$.
 Set $V:=\mathrm{span}\{f,g\}$ and $V^{\sigma}:=J(V)=\mathrm{span}\{Jf,Jg\}$.
-Then, $V^{\sigma}$ is invariant under $L(y)$ for every $y\in S$.
+Then $V^{\sigma}$ is invariant under $L(y)$ for every $y\in S$.
 Equivalently, $V$ is invariant under $R(\sigma(y))$ for every $y\in S$.
@@
 Thus $L(y)(Jf)\in V^{\sigma}$.
-We now show that $L(y)(Jg)\in V^{\sigma}$ for all $y\in S$. Indeed, choose $y_{0}\in S$ such that $h_{2}(y_{0})\neq 0$ (exists since $h_{1},h_{2}$ are linearly independent). From
+We now show that $L(y)(Jg)\in V^{\sigma}$ for all $y\in S$. Indeed, choose $y_{0}\in S$ such that $h_{2}(y_{0})\neq 0$ (this exists because $h_{2}\not\equiv 0$). From
 $L(y_{0})(Jf)=h_{1}(y_{0})Jf+h_{2}(y_{0})Jg$
@@
 We already know $L(z)(Jf)\in V^{\sigma}$ for all $z\in S$, so the right-hand side lies in $V^{\sigma}$. This proves the claim, hence $V^{\sigma}$ is $L$-invariant. ∎
+Remark 3.3 (Sharpness of the hypothesis $h_{2}\not\equiv 0$).
+Theorem 3.2 does not require linear independence of $h_{1}$ and $h_{2}$; in particular, it applies when $h_{1}=0$ and $h_{2}\not\equiv 0$.
+The proof uses only the existence of a point $y_{0}\in S$ with $h_{2}(y_{0})\neq 0$.
+This hypothesis is sharp.
+If $h_{2}\equiv 0$, the conclusion can fail.
+For example, let $S=C_{3}=\{e,a,a^{2}\}$, let $\sigma=\mathrm{id}_{S}$, let $f\equiv 1$, $h_{1}\equiv 1$, $h_{2}\equiv 0$, and let $g=\mathbf{1}_{\{e\}}$.
+Then
+$f(x\sigma(y))=1=f(x)h_{1}(y)+g(x)h_{2}(y)\qquad(\forall x,y\in S),$
+and $f,g$ are linearly independent. However,
+$L(a)g=\mathbf{1}_{\{a^{2}\}}\notin\mathrm{span}\{1,\mathbf{1}_{\{e\}}\}=\mathrm{span}\{f,g\}.$
+Hence $V^{\sigma}=V$ is not invariant under left translations.
 4. The Generalized Sine Law with an Anti-Automorphism
@@
 $f(xy)=f\bigl(x\sigma(\sigma(y))\bigr)=f(x)h_{1}(\sigma(y))+g(x)h_{2}(\sigma(y))=f(x)\widehat{h}_{1}(y)+g(x)\widehat{h}_{2}(y).$
-Lemma 4.3 (Central-element criterion for $\sigma$-conjugate span).
-Let $S$ be a semigroup and let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy
+Lemma 4.3 (Bridge criterion for $\sigma$-conjugate span).
+Let $S$ be a semigroup and let $\sigma:S\to S$ be an involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy
 $f(x\sigma(y))=f(x)\,h(y)+u(x)\,f(y)\qquad(\forall x,y\in S),$ (4.2)
@@
 $f\circ\sigma=\varepsilon f\qquad\text{for some }\varepsilon\in\{\pm 1\}.$ (4.3)
-If there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$, then
+If there exists $d\in S$ such that
+$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,u\},$ (4.4)
+then
 $u\circ\sigma\in\mathrm{span}\{f,u\}.$
-In particular, the conclusion holds whenever $S$ is commutative and $f\not\equiv 0$.
 Proof.
@@
 $f(\sigma(x)\sigma(y))=\varepsilon f(x)h(y)+(u\circ\sigma)(x)\,f(y).$
-Multiplying by $\varepsilon$ and using $\varepsilon^{2}=1$, we get the identity of functions
+Multiplying by $\varepsilon$ and using $\varepsilon^{2}=1$, we get
 $L(y)f=h(y)\,f+\varepsilon f(y)\,(u\circ\sigma)\qquad(\forall y\in S).$ (4.5)
-Now let $c\in Z(S)$ with $f(c)\neq 0$. Since $\sigma$ is surjective, pick $y_{0}\in S$ such that $\sigma(y_{0})=c$. Then
-$(R(\sigma(y_{0}))f)(x)=f(x\sigma(y_{0}))=f(xc)=f(cx)=(L(c)f)(x),$
-where we used $c\in Z(S)$. Applying Theorem 3.2 to (4.2) shows that $V:=\mathrm{span}\{f,u\}$ is invariant under $R(\sigma(\cdot))$, hence $R(\sigma(y_{0}))f\in V$, so $L(c)f\in V$.
-Finally, take $y=c$ in (4.4):
-$L(c)f=h(c)\,f+\varepsilon f(c)\,(u\circ\sigma).$
-Since $L(c)f\in V$ and $f(c)\neq 0$, we solve for $u\circ\sigma$:
-$u\circ\sigma=\frac{1}{\varepsilon f(c)}\,L(c)f-\frac{h(c)}{\varepsilon f(c)}\,f\in V.$
+Set $V:=\mathrm{span}\{f,u\}$. Taking $y=d$ in (4.5) and using (4.4), we obtain
+$L(d)f=h(d)\,f+\varepsilon f(d)\,(u\circ\sigma)\in V.$
+Since $f(d)\neq 0$, we can solve for $u\circ\sigma$:
+$u\circ\sigma=\frac{1}{\varepsilon f(d)}\bigl(L(d)f-h(d)f\bigr)\in V.$
 Thus $u\circ\sigma\in\mathrm{span}\{f,u\}$. ∎
+Corollary 4.4 (Central-element criterion).
+Let $S$ be a semigroup and let $\sigma:S\to S$ be an involutive anti-automorphism.
+Assume that $f,u,h:S\to F$ satisfy (4.2), where $f$ and $u$ are linearly independent, and that (4.3) holds for some $\varepsilon\in\{\pm 1\}$.
+If there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$, then
+$u\circ\sigma\in\mathrm{span}\{f,u\}.$
+In particular, the conclusion holds whenever $S$ is commutative and $f\not\equiv 0$.
+Proof.
+Set $V:=\mathrm{span}\{f,u\}$ and let $y_{0}:=\sigma(c)$. Then $\sigma(y_{0})=c$, so by (4.2),
+$R(c)f=R(\sigma(y_{0}))f=h(y_{0})\,f+f(y_{0})\,u\in V.$
+Since $c\in Z(S)$, we have $L(c)f=R(c)f\in V$. Thus (4.4) holds with $d=c$, and Lemma 4.3 gives the desired conclusion.
+For the commutative case, choose $d\in S$ with $f(d)\neq 0$. Since $\sigma$ is bijective, $d=\sigma(\sigma(d))$, and (4.2) with $y=\sigma(d)$ gives
+$R(d)f=R(\sigma(\sigma(d)))f\in V.$
+Because $S$ is commutative, $L(d)f=R(d)f\in V$, so (4.4) holds and Lemma 4.3 applies. ∎
+Remark 4.5.
+Strictly speaking, the algebraic derivations in Theorem 3.2, Lemma 4.3, and Corollary 4.4 do not rely on the linear independence of the functions involved; their conclusions remain valid even if the function pairs are linearly dependent.
+We state them with the linear independence hypothesis primarily for contextual consistency, as it seamlessly sets up the non-degenerate framework required for our main structural classification in the following important theorem.
 Theorem 4.6.
-Let $\sigma:S\to S$ be a surjective involutive anti-automorphism, and assume that there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$.
-Suppose $f,g:S\to F$ are linearly independent and satisfy (4.1).
+Let $\sigma:S\to S$ be an involutive anti-automorphism. Suppose $f,g:S\to F$ are linearly independent and satisfy (4.1). Assume that there exists $d\in S$ such that
+$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g\}.$ (4.6)
   1. (i)
@@
 $f(x\sigma(y))=f(x)\,h(y)+u(x)\,f(y),\qquad h=g+\gamma f,\ \ u=-g.$
-From Step 1 we already have $f\circ\sigma=-f$, i.e. (4.3) holds with $\varepsilon=-1$. Applying Lemma 4.3 to (4.2), we conclude that
+From Step 1 we already have $f\circ\sigma=-f$, i.e. (4.3) holds with $\varepsilon=-1$. Since u=gu=-g, the bridge hypothesis (4.6) reads
+$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,u\}.$
+Applying Lemma 4.3 to (4.2), we conclude that
 $u\circ\sigma\in\mathrm{span}\{f,u\}.$
@@
 Assume $\lambda=-1$. Then $g\circ\sigma=-g+\mu f$. Replacing $y$ by $\sigma(y)$ in (4.7) and using $f\circ\sigma=-f$ gives
-$f(xy)=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))=f(x)\bigl(-g(y)+\mu f(y)-\gamma f(y)\bigr)+g(x)f(y).$
+$f(xy)=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))=f(x)\bigl(-g(y)+(\mu-\gamma)f(y)\bigr)+g(x)f(y).$
 Let $\widehat{g}:=-g$. Then
@@
 $f(xy)$ $=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))$
-$=f(x)\bigl((g(y)+af(y))+\gamma(-f(y))\bigr)-g(x)(-f(y))$
 $=f(x)\bigl(g(y)+(a-\gamma)f(y)\bigr)+g(x)f(y).$
@@
 $-f(xy)$ $=f(\sigma(y)\sigma(x))=f(\sigma(y))\bigl(g(x)+\gamma f(x)\bigr)-g(\sigma(y))f(x)$
-$=(-f(y))g(x)+(-f(y))\gamma f(x)-\bigl(g(y)+af(y)\bigr)f(x)$
 $=-f(y)g(x)-g(y)f(x)-(a+\gamma)f(x)f(y).$
-Comparing the two expressions for f(xy)-f(xy) yields $\gamma=0$. With $\gamma=0$, (4.15) becomes
+Comparing the two expressions for f(xy)-f(xy), we obtain
+$2\gamma\,f(x)f(y)=0\qquad(\forall x,y\in S).$
+Since $f\not\equiv 0$, choose $x_{0}\in S$ with $f(x_{0})\neq 0$; taking $x=y=x_{0}$ and using $\operatorname{char}(F)\neq 2$, we get $\gamma=0$.
+With $\gamma=0$, (4.17) becomes
 f(xy)=f(x)g(y)+g(x)f(y)+af(x)f(y),f(xy)=f(x)g(y)+g(x)f(y)+a\,f(x)f(y),
@@
 $Q(z)=g_{1}(\sigma(z)),\qquad f(z)=\beta\,f(\sigma(z)).$ (4.24)
-Applying $\sigma$ to $f(z)=\beta f(\sigma(z))$ and using $\sigma^{2}={\rm id}$ gives $\beta^{2}=1$. Since $\beta\neq-1$, we conclude
+Applying $\sigma$ to $f(z)=\beta f(\sigma(z))$ and using $\sigma^{2}={\rm id}$, we obtain
+$f(\sigma(z))=\beta f(z)\qquad(\forall z\in S).$
+Hence
+$f(z)=\beta f(\sigma(z))=\beta^{2}f(z)\qquad(\forall z\in S).$
+Thus $(1-\beta^{2})f(z)=0$ for all $z\in S$. Since $f\not\equiv 0$, choose $z_{0}\in S$ with $f(z_{0})\neq 0$. Then $(1-\beta^{2})f(z_{0})=0$, hence $\beta^{2}=1$. Since $\beta\neq-1$, we conclude
 $\beta=1\quad\text{and hence}\quad f\circ\sigma=f.$ (4.25)
@@
 $f(x\sigma(y))=f(x)g_{1}(y)+g_{1}(x)f(y)\qquad(\forall x,y\in S).$
-This is exactly (4.2) with $u=h=g_{1}$ and $\varepsilon=1$ in (4.3). Applying Lemma 4.3, we get
+This is exactly (4.2) with $u=h=g_{1}$ and $\varepsilon=1$ in (4.3). Since $\mathrm{span}\{f,g_{1}\}=\mathrm{span}\{f,g\}$, the bridge hypothesis (4.6) gives
+$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g_{1}\}.$
+Applying Lemma 4.3, we get
 $g_{1}\circ\sigma\in\mathrm{span}\{f,g_{1}\}.$
@@
 hence $g\circ\sigma=g$. This proves (ii). ∎
-5. Examples: Concrete Illustrations of the Conjugation Identity
-We illustrate the conjugation identity on familiar groups (hence semigroups). In each example, $\sigma$ is an involutive anti-automorphism and $J$ denotes composition with $\sigma$,
+Corollary 4.7.
+The conclusions of Theorem 4.6 hold in particular if there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$.
+Proof.
+By (4.1), taking $y=\sigma(c)$ shows that
+$R(c)f=R(\sigma(\sigma(c)))f\in\mathrm{span}\{f,g\}.$
+Since $c\in Z(S)$, we have $L(c)f=R(c)f\in\mathrm{span}\{f,g\}$. Hence (4.6) holds with $d=c$, and Theorem 4.6 applies. ∎
+Remark 4.8.
+In the proof of Theorem 4.6, the dichotomy $\beta\in\{\pm 1\}$ and the parity relation $f\circ\sigma=\beta f$ are obtained before the bridge hypothesis (4.6) is used.
+The bridge assumption enters only through Lemma 4.3, which is invoked to show that $g\circ\sigma$ (or $g_{1}\circ\sigma$ in the case $\beta\neq-1$) belongs to the span of the basic solution pair.
+This span-closure step is then used to derive the standard $xy$-addition law and the explicit $\sigma$-transformation rule for $g$.
+The central-element condition of Corollary 4.7 is merely a convenient sufficient condition for (4.6); Example 5.4 in the sequel shows that it is strictly stronger.
+5. Examples and Further Illustrations
+We first illustrate the conjugation identity on familiar groups (hence semigroups).
+In each of the first three examples, $\sigma$ is an involutive anti-automorphism and $J$ denotes composition with $\sigma$,
 $(Jh)(x)=h(\sigma(x)).$
@@
 For the chosen finite-dimensional space $V$ of coordinate functions, we also describe explicitly how the left-translation operators $L(y)$ act on $V$.
-We emphasize that these examples are included solely to illustrate the conjugation identity $J\,R(\sigma(y))\,J=L(y)$ and the resulting left-translation action; they are independent of the additional centrality assumption used later in Theorem 4.4.
-5.1. Example 1: $GL_{n}(K)$ with $\sigma(A)=A^{\top}$ and column-coordinate functions
+We emphasize that Examples 1–3 below are included solely to illustrate the conjugation identity $J\,R(\sigma(y))\,J=L(y)$ and the resulting left-translation action (independent of the bridge hypothesis used in Theorem 4.6), whereas Example 4 focuses on this bridge hypothesis to show that it may hold even when no central element $c$ satisfies $f(c)\neq 0$.
+5.1. Example 1: GL(n,K) with transposition and column-coordinate functions
 Let $K$ be a field and let $S=GL_{n}(K)$ with $\sigma(A)=A^{\top}$.
+In this example, we take the coefficient field to be $F=K$.
 Then $\sigma$ is an involutive anti-automorphism since $(AB)^{\top}=B^{\top}A^{\top}$ and $(A^{\top})^{\top}=A$.
@@
 In particular, $L(y)V\subseteq V$ for every $y\in GL_{n}(K)$.
-5.2. Example 2: The symmetric group $S_{n}$ with $\sigma(\pi)=\pi^{-1}$
+5.2. Example 2: The symmetric group Sn
 Let $S=S_{n}$ and $\sigma(\pi)=\pi^{-1}$. Fix $j_{0}\in\{1,\dots,n\}$ and define, for $i=1,\dots,n$,
@@
 Thus $L(y)$ permutes the spanning set $\{f_{1},\dots,f_{n}\}$, and in particular $L(y)V\subseteq V$.
-5.3. Example 3: $SO(3,\mathbb{R})$ with $\sigma(g)=g^{-1}=g^{\top}$ and column functions
-Let $S=SO(3,\mathbb{R})$ and $\sigma(g)=g^{-1}=g^{\top}$. Write $g=(g_{ij})_{1\leq i,j\leq 3}$ and define
-$f_{1}(g):=g_{11},\qquad f_{2}(g):=g_{21},\qquad f_{3}(g):=g_{31},\qquad V:=\mathrm{span}\{f_{1},f_{2},f_{3}\}.$
-Verification of $JR(\sigma(y))J=L(y)$. We have $(Jf_{i})(X)=f_{i}(X^{\top})=X_{1i}$. For $y\in SO(3,\mathbb{R})$ and $x\in SO(3,\mathbb{R})$,
+5.3. Example 3: SO(3,R) and column functions
+Let $S=SO(3,\mathbb{R})$ and $\sigma(g)=g^{-1}=g^{\top}$. Here the coefficient field is $F=\mathbb{R}$. Write $g=(g_{i,\,j})_{1\leq i,\,j\leq 3}$ and define
+$f_{1}(g):=g_{1,\,1},\qquad f_{2}(g):=g_{2,\,1},\qquad f_{3}(g):=g_{3,\,1},\qquad V:=\mathrm{span}\{f_{1},f_{2},f_{3}\}.$
+Verification of $JR(\sigma(y))J=L(y)$. We have $(Jf_{i})(X)=f_{i}(X^{\top})=(X^{\top})_{i,\,1}=X_{1,\,i}$. For $y\in SO(3,\mathbb{R})$ and $x\in SO(3,\mathbb{R})$,
 $\big(JR(\sigma(y))Jf_{i}\big)(x)$ $=\big(R(\sigma(y))Jf_{i}\big)(\sigma(x))=\big(R(y^{\top})Jf_{i}\big)(x^{\top})$
@@
 Hence $JR(\sigma(y))J=L(y)$.
-Left action on $V$. Since $(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,x_{j1}$, we get for $y,x\in SO(3,\mathbb{R})$,
-$(L(y)f_{i})(x)=f_{i}(yx)=(yx)_{i1}=\sum_{j=1}^{3}y_{ij}\,f_{j}(x).$
+Left action on $V$. Since $(yx)_{i,\,1}=\sum_{j=1}^{3}y_{i,\,j}\,x_{j,\,1}$, we get for $y,x\in SO(3,\mathbb{R})$,
+$(L(y)f_{i})(x)=f_{i}(yx)=(yx)_{i,\,1}=\sum_{j=1}^{3}y_{i,\,j}\,f_{j}(x).$
 In particular, $L(y)V\subseteq V$.
@@
 Example 3. If $c_{1}f_{1}+c_{2}f_{2}+c_{3}f_{3}\equiv 0$, evaluate at $I$ to get $c_{1}=0$. Choose rotations sending $e_{1}$ to $e_{2}$ and $e_{1}$ to $e_{3}$ to get $c_{2}=c_{3}=0$.
-The examples above show concretely how Lemma 3.1 converts the family $R(\sigma(y))$ into left translations $L(y)$ via conjugation by $J$, and how this mechanism can be checked explicitly on natural finite-dimensional spaces of coordinate functions.
+5.4. Example 4: A non-central bridge instance on $S_{3}$
+Let $S=S_{3}$ and $\sigma(\pi)=\pi^{-1}$. Define
+$g(\pi):=\mathbf{1}_{A_{3}}(\pi),\qquad f(\pi):=\mathbf{1}_{S_{3}\setminus A_{3}}(\pi),$
+where $A_{3}$ denotes the alternating subgroup. Then $f$ and $g$ are linearly independent, and for all $x,y\in S_{3}$ one has
+$f(x\sigma(y))=f(xy^{-1})=f(x)g(y)+g(x)f(y).$
+Indeed, $xy^{-1}$ is odd if and only if $x$ and $y$ have opposite parity. Thus (4.1) holds with $\beta=1$ and $\gamma=0$.
+Choose any transposition $d\in S_{3}$. Then $f(d)=1$, and left multiplication by an odd permutation reverses parity, so
+$(L(d)f)(x)=f(dx)=g(x)\qquad(\forall x\in S_{3}).$
+Hence
+$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g\},$
+so the bridge hypothesis (4.6) holds.
+On the other hand, $Z(S_{3})=\{e\}$ and $f(e)=0$.
+Therefore no central element $c$ satisfies $f(c)\neq 0$.
+This shows that the bridge hypothesis of Theorem 4.6 is strictly weaker than the central-element condition in Corollary 4.7.
+The examples above show concretely how Lemma 3.1 converts the family $R(\sigma(y))$ into left translations $L(y)$ via conjugation by $J$, how this mechanism can be checked explicitly on natural finite-dimensional spaces of coordinate functions, and why the intrinsic bridge hypothesis in Theorem 4.6 is the right level of generality for the structural argument.
 6. Concluding Remarks
-The conjugation identity $J\,R(\sigma(y))\,J=L(y)$ provides the key link needed to transfer the anti-automorphic right-translate action into left translations, thereby avoiding the order-reversal obstruction in the family $R(\sigma(\cdot))$.
-This left-translation viewpoint yields a clean Levi–Civita closure principle in the anti-automorphic setting and leads to analogues of Stetkær’s structural results for the generalized sine law, including the dichotomy $\beta\in\{\pm 1\}$ and the induced $\sigma$-parity rule for $f$.
-Moreover, under a mild additional centrality hypothesis (e.g.
-for commutative semigroups), one also recovers the full $\sigma$-transformation rule for $g$.
+The conjugation identity $J\,R(\sigma(y))\,J=L(y)$ provides an intrinsic operator interpretation of the order-reversing family $R(\sigma(\cdot))$ that arises from an involutive anti-automorphism.
+Although an equation of the form $f(x\sigma(y))=\cdots$ can also be rewritten by the substitution $y\mapsto\sigma(y)$, the formula $J\,R(\sigma(y))\,J=L(y)$ shows directly how the anti-automorphic parameter family is converted into a genuine left-translation action.
+This left-translation viewpoint yields a natural Levi–Civita closure principle in the anti-automorphic setting and leads, under the bridge hypothesis used in Theorem 4.6, to analogues of Stetkær’s structural results for the generalized sine law, including the dichotomy $\beta\in\{\pm 1\}$, the standard $xy$-addition law, and the expected $\sigma$-transformation rules.
+In the proof, the parity rule $f\circ\sigma=\beta f$ is obtained before the bridge step, whereas the bridge hypothesis is used to control $g\circ\sigma$ and thereby to pass from the $x\sigma(y)$-law to the standard $xy$-addition law.
+Corollary 4.7 shows that the existence of a central element $c$ with $f(c)\neq 0$ is a convenient sufficient condition for the bridge hypothesis, while Example 5.4 shows that this sufficient condition is not necessary.
 The examples illustrate the conjugation mechanism concretely on standard semigroups and show how the relevant finite-dimensional spaces of coordinate functions behave under left translations.
+Acknowledgment
+The author is deeply grateful to the handling editor for the careful handling of the manuscript and to the anonymous referee for reading the paper thoroughly and for providing valuable comments and suggestions, which led to a significant improvement in the presentation.
 Funding.